OpenStudy (anonymous):

The ceiling in your house is 9 feet. The floor of a room you want to paint has dimensions of 12' by 16'/ The coverage of the paint will you need to buy if you want to double coat the walls and the ceiling? What will the material cost be, if paint is on sale for $14.95 per gallon down at the local hardware store? How much paint will be left over, if paint is sold only by the gallon?

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

@dan815 can u help?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: we have to paint the ceiling plus four walls, namely the lateral area plus the ceiling

OpenStudy (michele_laino):

now, the area of the ceiling is: 12*16=...?

OpenStudy (michele_laino):

whereas the area of four walls is given by the subsequent computation: base perimeter * height of room= =2*(12+16)*9=...? @katlin95

OpenStudy (anonymous):

omg! sorry i didnt get back to u sooner!!

OpenStudy (anonymous):

504

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

ok! the total surface is: 504+192=...?

OpenStudy (anonymous):

523

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now, we have to double that quantity, since we have to double coat the ceiling and the walls

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

so how would we go about doing that?

OpenStudy (michele_laino):

the requested quantity, is: 2*(504+192)=2*696=...?

OpenStudy (michele_laino):

sorry 523 is wong since 504+192=696

OpenStudy (michele_laino):

wrong*

OpenStudy (anonymous):

wait, im confused lol

OpenStudy (michele_laino):

the surface ceiling plus walls is 12*16+504=696 is it ok?

OpenStudy (anonymous):

well, my answer book shows an answer of "need 4 gallons, 0.91 gallons left over

OpenStudy (michele_laino):

how much surface we can paint using 1 gallon?

OpenStudy (michele_laino):

namely, we know the area which we have to paint, nevertheless we don't know how many gallons we need

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

and I don't know how to compute that number of gallons. Maybe some data are missing in your problem?

OpenStudy (anonymous):

ya, the question seems confusing. Ill just move onto the next question:)

OpenStudy (anonymous):

I have one big question that I could really use ur help on, coul u help me?

OpenStudy (michele_laino):

ok!